Two objects attract each other with a gravitational force of magnitude 1.01 x 10-8 N when separated by 20.2 cm. If the total mass of the two objects is 5.11 kg, what is the mass of each? heavier mass lighter mass kg kg Need Help?Watch It
Answer #1

The force of attraction is

\frac{GMm}{r^2}=\frac{6.67\times 10^{-11}Mm}{0.202^2}=1.01\times 10^{-8}

This gives

Mm=6.179

and the total mass is

M+m=5.11\,kg

Solving the above two equation gives

M=3.146\,\,kg (heavier mass)

and

m=1.964\,\,kg (lighter mass)